OpenStudy (anonymous):

A triangle has a base of 12 meters and a height of 9 meters. What is the area of the triangle? A. 21 square meters B. 42 square meters C. 54 square meters D. 108 square meters

OpenStudy (harsha19111999):

c option area=1/2 base height

hartnn (hartnn):

\(Area ~ of ~ triangle ~ A= (1/2) base \times height \)

OpenStudy (harsha19111999):

So 1/2*12*9=54

OpenStudy (insa):

54

OpenStudy (mathmale):

This formula for the area of a triangle is very, very definitely one that you should know by heart. Did you look up "area of a triangle" in your online study materials, do an Internet search for this term, or consult a textbook the way we old timers used to?